OpenStudy (anonymous):

A solid has 6 faces and 8 vertices. How many edges does it have?

OpenStudy (nurali):

number of faces + number of vertices minus the number of edges = 2. It can be written like this : F + V - E = 2 lets plug in our numbers... 6 + 8 - E = 2 14- E = 2 14 - 2 = E 12= E

OpenStudy (nurali):

Welcome to Openstudy.